A dres was marked dow at 30% .the dress now costs ph280.00.what was the price before the mark down?
1
Expert's answer
2013-04-18T10:48:05-0400
Let x is a price before mark down.
After the mark down the price is x - 0.3x = 0.7x and the price is ph280.
Then we can make the equation:
0.7x=280
x=280/0.7=400
Answer: the price before the mark down was ph400.00